OpenStudy (anonymous):

A teaspoon equals 1/6 of a fluid ounce. A cup equals 8 fluid ounces. How many teaspoons are there in 1 cup? Show you steps. = )

OpenStudy (dumbcow):

A:1 cup = 8 oz 1 tsp = 1/6 oz multiply 6 on both sides 6 tsp = 1 oz substitute 6 tsp in for 1 oz in A -> 1 cup = 8*6tsp 1 cup = 48 tsp

OpenStudy (anonymous):

Thank You = )